# Interpret the estimand framework from a causal inference perspective

### v1

@@ Abstract
The estimand framework proposed by ICH in 2017 has brought fundamental changes in the pharmaceutical industry. It clearly describes how a treatment effect in a clinical question should be precisely defined and estimated, through attributes including treatments, endpoints and intercurrent events. However, ideas around the estimand framework are commonly in text, and different interpretations on this framework may exist. This article aims to interpret the estimand framework through its underlying theories, the causal inference framework based on potential outcomes. The statistical origin and formula of an estimand is given through the causal inference framework, with all attributes translated into statistical terms. How five strategies proposed by ICH to analyze intercurrent events are incorporated in the statistical formula of an estimand is described, and a new strategy to analyze intercurrent events is also suggested. The roles of target populations and analysis sets in the estimand framework are compared and discussed based on the statistical formula of an estimand. This article recommends continuing study of causal inference theories behind the estimand framework and improving the estimand framework with greater methodological comprehensibility and availability.
@@ 1 Introduction
The estimand framework was drafted in 2017 by ICH in Efficacy Guideline E9(R1) as addendum to Efficacy Guideline E9 and later published in 2019 [2, 3]. It has gained increasing attention in the pharmaceutical industry. Many clinical trials have used the estimand framework to develop new drugs for both oncology and non-oncology diseases [4], and professional working groups such as the Oncology Estimand Working Group (https://oncoestimand.github.io/oncowg_webpage/docs/) have been initiated to study how the estimand framework should be incorporated in pharmaceutical practices.
The estimand framework aims to clearly define a clinical question. This means to clearly define a treatment effect in this clinical question. The estimand framework introduces and compares estimands, estimators and estimates with regard to statistical roles in estimation of treatment effects. An estimand is a precise definition of a treatment effect in a clinical question, an estimator is a statistical method that estimates this estimand, and an estimate is a result from this estimator [3]. Five attributes are proposed for an estimand. They are treatments, endpoints, a target population, intercurrent events, population-level summary [3]. Treatments are medical interventions that patients would take in clinical trials. Endpoints are outcomes used to assess efficacy and safety of treatments. A target population is a group of patients with medical conditions of clinical interest. Intercurrent events are events that happen after treatment initiation and affect the definition of a treatment effect. Population-level summary is an approach to estimate the treatment effect in a target population.
Intercurrent events are frequent in practice but conceptually novel. E9(R1) listed many examples for intercurrent events, such as use of concomitant therapies, treatment switching and death before endpoint measurement. To suit different study objectives, five strategies have been proposed to reduce bias from intercurrent events. They are the treatment policy strategy, the hypothetical strategy, the composite variable strategy, the while on treatment strategy, the principal stratum strategy [3].
@@ 2 The causal inference framework for estimands
A causal inference framework is based on the potential outcome framework [1, 5]. Suppose there is an ideal two-arm randomized controlled clinical trial, with full compliance to treatment and no intercurrent events. This trial has a sample size of $N$, a binary treatment $X$, a continuous endpoint $Y$, a randomization scheme $R$ and some confounders $C$ that affect both $X$ and $Y$. $X$, $Y$, $R$ are random vectors of length $N$, and $C$ is a random matrix of row dimension $N$. $X_{i}$, $Y_{i}$, $R_{i}$, $C_{i}$ represent random variables or vectors for a participant $i$, where $i\in\{1,2,3,\ldots,N\}$. $R_{i}=0$ means that the participant is assigned to the control arm and $R_{i}=1$ means that the participant is assigned to the treatment arm. $X_{i}(R_{i}=0)=0$ means that the participant would take the control treatment if assigned to the control arm and $X_{i}(R_{i}=1)=1$ means that the participant would take the experimental treatment that is of primary clinical interest if assigned to the treatment arm. $Y_{i}(X_{i}(R_{i}=0)=0)$ would be the endpoint if the participant took the control treatment and $Y_{i}(X_{i}(R_{i}=1)=1)$ would be the endpoint if the participant took the experimental treatment. $R_{i}$, $X_{i}$ and $Y_{i}$ are potential outcomes. They are not real because we hypothesize what would happen if the participant took the control treatment or instead took the experimental treatment. Figure 1 is a causal directed acyclic graph that shows the causal relationships between $X$, $Y$, $R$, $C$. $R$ affects $Y$ only through $X$.
@@
A clinical question is to estimate the overall treatment effect of taking the experimental treatment versus taking the control treatment on the endpoint over all participants. For the participant $i$, if he had $X_{i}(R_{i}=0)=0$ then he would have $Y_{i}(X_{i}(R_{i}=0)=0)$, and if had $X_{i}(R_{i}=1)=1$ then he would have $Y_{i}(X_{i}(R_{i}=1)=1)$. Hence, for this participant, an individual treatment effect is $Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0)$. This individual treatment effect controls confounders on the endpoint within the same participant and means how the endpoint would change when only the treatment condition changes. There are $N$ individual treatment effects. They can be assumed to be equal or unequal. Generally, we are interested in an average treatment effect (ATE) among all participants. The ATE is an average of all individual treatment effects and can be described as $E(Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0)~{}|~{}C)$. And the estimand in this clinical question is the ATE. If there is no confounder, the ATE would become $E(Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0))$. Without adjusting for existing confounding effects, the ATE estimation may be biased.
Further, $E(Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0)~{}|~{}C)=E(Y_{i}(X_{i}(R_{i}=1)=1)~{}|~{}C)-E(Y_{i}(X_{i}(R_{i}=0)=0)~{}|~{}C)$. This also means that the estimand is the difference between the ATE of taking the experimental treatment among all participants and the ATE of taking the control treatment among all participants. The problem is that, in real world, each participant only takes one kind of treatment, and only one of the two hypothetical situations with regard to two arms can happen. In this ideal trial, the observed randomized assignment $R_{i}^{o}$ is either 0 and 1, the observed treatment $X_{i}^{o}$ is either $X_{i}(R_{i}=0)$ or $X_{i}(R_{i}=1)$, and the observed outcome $Y_{i}^{o}(X_{i}^{o})$ is either $Y_{i}(X_{i}(R_{i}=0))$ or $Y_{i}(X_{i}(R_{i}=1))$. Hence, the relationship between potential outcomes and observed variables can be described as
@@
What we can estimate from the actual trial data is the difference between the ATE from participants who take the experimental treatment in the treatment arm and the ATE from participants who take the control treatment in the control arm, that is, $E(Y^{o}~{}|~{}X^{o}=1,R^{o}=1,C)-E(Y^{o}~{}|~{}X^{o}=0,R^{o}=0,C)$. With certain assumptions including the stable unit treatment value assumption, the randomization assumption and the exclusion restriction assumption, it can proved that $E(Y^{o}~{}|~{}X^{o}=1,R^{o}=1,C)-E(Y^{o}~{}|~{}X^{o}=0,R^{o}=0,C)=E(Y^{o}~{}|~{}X^{o}=1,C)-E(Y^{o}~{}|~{}X^{o}=0,C)=E(Y(X_{i}(R_{i}=1)=1)~{}|~{}C)-E(Y(X_{i}(R_{i}=0)=0)~{}|~{}C)$. Without these assumptions, this equality chain will not hold.
@@
$Y^{o}=\beta_{0}+\beta_{1}X^{o}+\beta_{2}C+\varepsilon,E(\varepsilon)=0,Var(\varepsilon)=\sigma^{2},$
where $E(Y^{o}~{}|~{}X^{o},C)=\beta_{0}+\beta_{1}X^{o}+\beta_{2}C$. Then, $E(Y^{o}~{}|~{}X^{o}=0,C)=\beta_{0}+\beta_{2}C$ and $E(Y^{o}~{}|~{}X^{o}=1,C)=\beta_{0}+\beta_{1}+\beta_{2}C$, which implies that $E(Y^{o}~{}|~{}X^{o}=1,C)-E(Y^{o}~{}|~{}X^{o}=0,C)=\beta_{1}$. Hence, $\beta_{1}$ equals to the ATE and thus is an estimator to the estimand from this linear model. After the linear model is built with the actual trial data, an estimate $\hat{\beta}_{1}$ on $\beta_{1}$ would be obtained, and $\hat{\beta}_{1}$ would be an estimate to the estimand.
The statistical formula of the estimand as $E(Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0)~{}|~{}C)$ has already contained five attributes proposed in the estimand framework. The attribute of treatments is represented by $X$. The attribute of endpoints is represented by $Y$. The attribute of population-level summary is represented by the difference between two ATEs. The attribute of intercurrent events is not necessary here since no intercurrent events are assumed. The attribute of a target population is implicitly stated as the trial population and can be made explicit in the definition of the estimand. We can introduce a selection variable $S$ that indicates how the target population is selected from the general public and update the estimand to $E(Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0)~{}|~{}C,S)$.
The causal inference framework is not restricted to the above situation. It is applicable to multivalued treatments, discrete outcomes and Bayesian methods. It is also the underlying theory that applies to any effort of estimating treatment effects. We may have used it, even if the causal inference framework does not appear explicitly in inference, because potential outcomes have been transformed to observed variables in statistical models and we only see observed variables. Causal inference is a mature, continuing research area in Statistics. Methods may be available for difficult practical problems. Hence, it would be recommended that we try to interpret the estimand from a causal inference perspective based on our study objectives, and understand or choose relevant methods to estimate estimands.
@@ 3 Causal inference with intercurrent events
The attribute of intercurrent events can also be incorporated in the statistical formula of an estimand. Following the trial example in the last section, suppose now the trial has intercurrent events shown in figure 2. Also suppose the endpoint is not related to death and the second assessment on the endpoint is used for primary analysis. We use these intercurrent events to discuss the statistical formula of an estimand in different strategies.
@@ 3.1 The treatment policy strategy
The treatment policy strategy is to include intercurrent events in the treatment definition. In case (1), when concomitant therapies are used, the endpoint may also be affected by concomitant therapies in addition to the control and experimental treatments. Hence, we cannot estimate the ATE of the experimental treatment versus the control treatment on the endpoint without adjusting for concomitant therapies. Through the treatment policy strategy, we re-define $X$ as the control or experimental treatment plus any concomitant therapy. Then we can use all participants from case (1) in data analysis. The estimand formula is still $E(Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0)~{}|~{}C)$, but now the estimand has been changed into the ATE of the experimental treatment plus any concomitant therapy versus the control treatment plus any concomitant therapy.
This strategy does not estimate the pure effect of the experimental treatment. When concomitant therapies are common for use with the experimental treatment in real world, this strategy would be appropriate.
@@ 3.2 The hypothetical strategy
In case (1), if concomitant therapies are prohibited by the protocol, we cannot include concomitant therapies in the treatment definition and the endpoint assessments after prohibited therapies are used are not appropriate for analysis. Through the hypothetical strategy, we make the second assessment on the endpoint after prohibited therapies are used missing and impute missing endpoint assessments as if there were only the control and experimental treatments. In case (2), the second assessment on the endpoint is naturally missing due to death. Through the hypothetical strategy, we impute missing endpoint assessments with the assumption that participants are alive at this time. For both cases, the estimand formula is still $E(Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0)~{}|~{}C)$, and the meaning of the estimand is not changed.
This strategy estimates the pure effect of the experimental treatment. When the pure treatment effect is of major clinical importance in real world, this strategy would be appropriate.
@@ 3.3 The composite variable strategy
The composite variable strategy is to include intercurrent events in the endpoint definition. In case (2), suppose the endpoint is a disability index measured by questionnaires and it does not consider death. Since participants from case (2) die by the second assessment, the questionnaire measurement is not done at the second assessment. However, death indicates no physical function. Through the composite variable strategy, death is included in the worst disability category and thus can be used in data analysis. The estimand formula is still $E(Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0)~{}|~{}C)$, but now the definition of $Y$ has been changed from disability measurement with no death included to disability measurement with death included in the worst disability category.
This strategy estimates the pure effect of the experimental treatment on a different endpoint. When composite endpoints are realistic in real world, this strategy would be appropriate.

### v2

@@ Abstract
The estimand framework proposed by ICH in 2017 has brought fundamental changes in the pharmaceutical industry. It clearly describes how a treatment effect in a clinical question should be precisely defined and estimated, through attributes including treatments, endpoints and intercurrent events. However, ideas around the estimand framework are commonly in text, and different interpretations on this framework may exist. This article aims to interpret the estimand framework through its underlying theories, the causal inference framework based on potential outcomes. The statistical origin and formula of an estimand is given through the causal inference framework, with all attributes translated into statistical terms. We describe how five strategies proposed by ICH to analyze intercurrent events are incorporated in the statistical formula of an estimand, and we also suggest a new strategy to analyze intercurrent events. The roles of target populations and analysis sets in the estimand framework are compared and discussed based on the statistical formula of an estimand. This article recommends continuing studying causal inference theories behind the estimand framework and improving the estimand framework with greater methodological comprehensibility and availability.
@@ 1 Introduction
The estimand framework was drafted in 2017 by ICH in Efficacy Guideline E9(R1) as addendum to Efficacy Guideline E9 and later published in 2019 [1, 2]. It has gained increasing attention in the pharmaceutical industry and the academia [4, 5, 3, 6, 7, 8]. Many clinical trials have used the estimand framework to develop new drugs for both oncology and non-oncology diseases, and professional working groups such as the Oncology Estimand Working Group (https://oncoestimand.github.io/oncowg_webpage/docs/) have been initiated to study how the estimand framework should be incorporated in pharmaceutical practices [9].
The estimand framework aims to clearly define a clinical question. This means to clearly define a treatment effect in this clinical question. The estimand framework introduces and compares estimands, estimators and estimates with regard to statistical roles in estimation of treatment effects. An estimand is a precise definition of a treatment effect in a clinical question, an estimator is a statistical method that estimates this estimand, and an estimate is a result from this estimator [2]. Five attributes are proposed for an estimand. They are treatments, endpoints, a target population, intercurrent events, and a population-level summary [2]. Generally, Treatments are medical interventions that patients would take in clinical trials. Endpoints are outcomes used to assess efficacy and safety of treatments. A target population is a group of patients with medical conditions of clinical interest. Intercurrent events are events that happen after treatment initiation and affect the definition of a treatment effect. A population-level summary is an approach to estimate the treatment effect in a target population.
Intercurrent events are frequent in practice but conceptually novel. E9(R1) listed many examples for intercurrent events, such as use of concomitant therapies, treatment switching and death before endpoint measurement. To suit different study objectives, five strategies have been proposed to reduce bias from intercurrent events. They are the treatment policy strategy, the hypothetical strategy, the composite variable strategy, the while on treatment strategy, and the principal stratum strategy [2].
@@ 2 The causal inference framework for estimands
A causal inference framework is based on the potential outcome framework [10, 11, 12]. Suppose there is an ideal two-arm randomized controlled clinical trial, with full compliance to treatment and no intercurrent events. This clinical trial has a sample size of $N$, a binary treatment $X$, a continuous endpoint $Y$, a randomization scheme $R$ and some confounders $C$ that affect both $X$ and $Y$. $X$, $Y$, $R$ are random vectors of length $N$, and $C$ is a random matrix of row dimension $N$. $X_{i}$, $Y_{i}$, $R_{i}$, $C_{i}$ represent random variables or vectors for the participant $i$, where $i\in\{1,2,3,\ldots,N\}$. $R_{i}=0$ means that the participant is assigned to the control arm, and $R_{i}=1$ means that the participant is assigned to the treatment arm. $X_{i}(R_{i}=0)=0$ means that the participant would take the control treatment if assigned to the control arm, and $X_{i}(R_{i}=1)=1$ means that the participant would take the experimental treatment that is of primary clinical interest if assigned to the treatment arm. $Y_{i}(X_{i}(R_{i}=0)=0)$ would be the endpoint if the participant took the control treatment as assigned, and $Y_{i}(X_{i}(R_{i}=1)=1)$ would be the endpoint if the participant took the experimental treatment as assigned. $R_{i}$, $X_{i}$ and $Y_{i}$ are potential outcomes. They are not real because we hypothesize what would happen if the participant took the control treatment or instead took the experimental treatment. Figure 1 is a causal directed acyclic graph that shows the causal relationships between $X$, $Y$, $R$, $C$. $R$ affects $Y$ only through $X$.
@@
A clinical question is to estimate the overall treatment effect of taking the experimental treatment versus taking the control treatment on the endpoint over all participants. For the participant $i$, if he had $X_{i}(R_{i}=0)=0$ then he would have $Y_{i}(X_{i}(R_{i}=0)=0)$, and if he had $X_{i}(R_{i}=1)=1$ then he would have $Y_{i}(X_{i}(R_{i}=1)=1)$. For this participant, the individual treatment effect (ITE) is the difference between two potential outcomes of $Y_{i}$. That is,
$ITE$ $=$ $Y_{i}(X_{i}(R_{i}=1)=1)-Y_{i}(X_{i}(R_{i}=0)=0).$
This individual treatment effect controls confounders on the endpoint within the same participant and means how the endpoint would change when only the treatment condition changes. There are $N$ individual treatment effects. They can be assumed to be equal or unequal. Suppose we are interested in an average treatment effect (ATE) among all participants. The ATE is an average of all individual treatment effects and can be described as
$ATE$ $=$ $E(Y(X(R=1)=1)-Y(X(R=0)=0)~|~C).$
In the clinical question, the estimand is the ATE. If there is no confounder, the ATE would become E(Y(X(R=1)=1)Y(X(R=0)=0))E(Y(X(R=1)=1)-Y(X(R=0)=0)). Without adjusting for existing confounding effects, the ATE estimation may be biased.
Further, based on expectation properties, we have
$ATE$ $=$ $E(Y(X(R=1)=1)~|~C)-E(Y(X(R=0)=0)~|~C).$
This also means that the estimand is the difference between the average treatment effect of taking the experimental treatment among all participants and the average treatment effect of taking the control treatment among all participants. The problem is that, in the real world, each participant only takes one kind of treatment, and only one of the two hypothetical situations with regard to two arms can happen. In this clinical trial setting, for each participant, the observed randomized assignment $R_{i}^{o}$ is either 0 and 1, the observed treatment $X_{i}^{o}$ is either $X_{i}(R_{i}=0)$ or $X_{i}(R_{i}=1)$, and the observed outcome $Y_{i}^{o}(X_{i}^{o})$ is either $Y_{i}(X_{i}(R_{i}=0))$ or $Y_{i}(X_{i}(R_{i}=1))$. The relationship between potential outcomes and observed variables can be described as
@@
What we can estimate from the actual clinical trial data is the difference (D) between the average treatment effect from participants who take the experimental treatment in the treatment arm and the average treatment effect from participants who take the control treatment in the control arm. That is,
$D$ $=$ $E(Y^{o}~|~X^{o}=1,R^{o}=1,C)-E(Y^{o}~|~X^{o}=0,R^{o}=0,C).$
With certain assumptions, including the stable unit treatment value assumption, the randomization assumption and the exclusion restriction assumption, it can be proved that D equals to the ATE, through
$D$ $=$ $E(Y^{o}~|~X^{o}=1,C)-E(Y^{o}~|~X^{o}=0,C),$ (2.1)
$=$ $E(Y(X(R=1)=1)~|~C)-E(Y(X(R=0)=0)~|~C).$
@@
$Y^{o}=\beta_{0}+\beta_{1}X^{o}+\beta_{2}C+\varepsilon,$
$E(\varepsilon)=0,Var(\varepsilon)=\sigma^{2},$
where $E(Y^{o}~|~X^{o},C)=\beta_{0}+\beta_{1}X^{o}+\beta_{2}C$. Then, we have two equations with different observed treatment levels as
$E(Y^{o}~|~X^{o}=0,C)$ $=$ $\beta_{0}+\beta_{2}C,$
$E(Y^{o}~|~X^{o}=1,C)$ $=$ $\beta_{0}+\beta_{1}+\beta_{2}C,$
which implies that
$\beta_{1}=E(Y^{o}~|~X^{o}=1,C)-E(Y^{o}~|~X^{o}=0,C),$ (2.2)
Hence, $\beta_{1}$ equals to the ATE through equations 2.1 and 2.2. It is an estimator to the estimand from this linear model. After the linear model is built with the actual clinical trial data, an estimate $\hat{\beta}_{1}$ on $\beta_{1}$ would be obtained. $\hat{\beta}_{1}$ is an estimate to the estimand.
The statistical formula of the estimand as E(Y(X(R=1)=1)Y(X(R=0)=0)|C)E(Y(X(R=1)=1)-Y(X(R=0)=0)~|~C) has already contained five attributes proposed in the estimand framework. The attribute of treatments is represented by $X$. The attribute of endpoints is represented by $Y$. The attribute of a population-level summary is represented by the difference in the average treatment effect between taking the experimental treatment and taking the control treatment among all participants. The attribute of intercurrent events is not necessary in this clinical trial setting because no intercurrent events are assumed. The attribute of a target population is implicitly stated as the clinical trial population and can be made explicit in the definition of the estimand. We can introduce a selection variable $S$ that indicates how the target population is selected from the general public and update the estimand to E(Y(X(R=1)=1)Y(X(R=0)=0)|C,S)E(Y(X(R=1)=1)-Y(X(R=0)=0)~|~C,S).
The causal inference framework is not restricted to the above situation. It is applicable to multivalued treatments, discrete outcomes, nonlinear risk comparisons and Bayesian methods. It is also the underlying theory that applies to any effort of estimating treatment effects. We may have used it, even if the causal inference framework does not appear explicitly in inference, because potential outcomes have been transformed to observed variables in statistical models and we only see observed variables. Causal inference is a mature, continuing research area in Statistics. Methods may be available for difficult practical problems. Hence, it would be recommended that we try to interpret the estimand from a causal inference perspective based on our study objectives, and understand or choose relevant methods to estimate estimands.
@@ 3 Causal inference with intercurrent events
The attribute of intercurrent events can also be incorporated in the statistical formula of an estimand. Following the clinical trial example in the last section, suppose now the clinical trial has intercurrent events as shown in figure 2. Also suppose the endpoint is not related to death and the second assessment on the endpoint is used for primary analysis. We use these intercurrent events to discuss the statistical formula of an estimand in different strategies.
@@ 3.1 The treatment policy strategy
The treatment policy strategy is to include intercurrent events in the treatment definition. In case (1), when concomitant therapies are used, the endpoint may also be affected by concomitant therapies in addition to the control and experimental treatments. Hence, we cannot estimate the ATE of the experimental treatment versus the control treatment on the endpoint without adjusting for concomitant therapies. Through the treatment policy strategy, we re-define $X$ as the control or experimental treatment plus any concomitant therapy. Then we can use all participants from case (1) in data analysis. The estimand formula is still E(Y(X(R=1)=1)Y(X(R=0)=0)|C)E(Y(X(R=1)=1)-Y(X(R=0)=0)~|~C), but now the estimand has been changed into the ATE of the experimental treatment plus any concomitant therapy versus the control treatment plus any concomitant therapy.
This strategy does not estimate the pure effect of the experimental treatment. When concomitant therapies are common for use with the experimental treatment in the real world, this strategy would be appropriate.
@@ 3.2 The hypothetical strategy
In case (1), if concomitant therapies are prohibited by the clinical trial protocol, we cannot include concomitant therapies in the treatment definition and the endpoint assessments after prohibited therapies are used are not appropriate for analysis. Through the hypothetical strategy, we make the second assessment on the endpoint after prohibited therapies are used missing and impute missing endpoint assessments as if there were only the control and experimental treatments. In case (2), the second assessment on the endpoint is naturally missing due to death. Through the hypothetical strategy, we impute missing endpoint assessments with the assumption that participants are alive at this time. For both cases, the estimand formula is still E(Y(X(R=1)=1)Y(X(R=0)=0)|C)E(Y(X(R=1)=1)-Y(X(R=0)=0)~|~C), and the meaning of the estimand is not changed.
This strategy estimates the pure effect of the experimental treatment. When the pure treatment effect is of major clinical importance in the real world, this strategy would be appropriate.
@@ 3.3 The composite variable strategy
The composite variable strategy is to include intercurrent events in the endpoint definition. In case (2), suppose the endpoint is a disability index measured by questionnaires and it does not consider death. Since participants die by the second assessment, the questionnaire measurement is not done at the second assessment. However, death already indicates no physical function. Through the composite variable strategy, death is included in the worst disability category and thus can be used in data analysis. The estimand formula is still E(Y(X(R=1)=1)Y(X(R=0)=0)|C)E(Y(X(R=1)=1)-Y(X(R=0)=0)~|~C), but now the definition of $Y$ has been changed from disability measurement with no death included to disability measurement with death included in the worst disability category.
This strategy estimates the pure effect of the experimental treatment on a different endpoint. When composite endpoints are realistic in the real world, this strategy would be appropriate.